# Gravitational lensing of galaxy clustering

### v1

@@ 1 Introduction
There is large scale structure (LSS) in the universe, as evidenced by the spatial clustering of galaxies (Peebles, 1980). This LSS deflects light as it propagates through the inhomogeneous universe (Bartelmann & Schneider, 2001). These deflections can occasionally lead to spectacular Einstein rings and multiple images, but usually will lead to only small distortions to the shapes and brightnesses of background objects. This has been measured in galaxy surveys through the changes in the detailed shapes of galaxies (Luo et al., 2018; Gatti et al., 2021), and has been measured in the distortion to the characteristic clustering patterns in the cosmic microwave background (Aghanim et al., 2020; Omori et al., 2023; Madhavacheril et al., 2023; Ade et al., 2015; BICEP/Keck Collaboration et al., 2022).
@@
In the case of CMB lensing, the workhorse tool that is widely used is a quadratic estimator Okamoto & Hu (2003), which amounts to multiplying together two CMB maps that each have their own filtering. The filtering of the maps is chosen to create an unbiased, minimum variance estimator, which optimizes the SNR of the resulting lens reconstruction. If this lens reconstruction is then correlated with a galaxy-based LSS tracer, the result could be expressed as a 3-pt function, as it involves products of two CMB maps and an LSS map. As we will show below, for the similar case of lensing of the galaxy clustering distribution by foreground matter, the reconstruction of the foreground lensing potential will involve two versions of the background galaxy number count field, and then cross-correlation with an LSS tracer such as cosmic shear can be expressed as a galaxy-galaxy-shear 3-pt function. If we instead used the foreground galaxy counts, then there is a 3-pt function that would be quadratic in the background galaxy number counts and linear in foreground galaxy number counts. In general, the three fields going into these estimates can be from diverse LSS tracers: the background 2-pt function could be cross-correlations between different LSS tracers, and the foreground could be taken to be any LSS tracer of the foreground matter. In this work, we only look at the simplest case of using the auto-correlation of galaxies in the background being lensed by foreground matter in a way that leads to a measured shear of the same background material, mainly to keep the accounting simple.
@@
An important point of focus in each of these studies has been the non-Gaussianity of the background field being lensed. It has been shown (Schaan et al., 2018; Foreman et al., 2018; Feng & Holder, 2019) that there can be substantial biases in the recovered lensing signal if the non-Gaussianity is ignored. Nevertheless, for this preliminary investigation we will disregard the non-Gaussianity of the field and its uncertainties in what follows, so the results presented below will no doubt require future work to study these effects. These corrections arise because non-linear gravitational collapse leads to an intrinsic background of spatial variation in the statistics of small-scale clustering, the magnitude of which depends on factors including the spatial scales probed, redshift, and survey parameters (Foreman et al., 2018). However, some lensing estimators that are optimal for a Gaussian source field have been shown to be resistant to non-Gaussian noise. Foreman et al. (2018) constructed a bias-hardened estimator that included contributions from the density field trispectrum that corrected for mild gravitational nonlinearities, while Lu et al. (2010) calibrated their estimator with N-body simulations to create a set of lensing weights that are optimal for non-Gaussian fields.
@@
In this work, we closely follow the procedures described by Schaan & Ferraro (2019), who present a method for lensing reconstruction that is less affected by foreground contamination and non-Gaussianities. In Schaan & Ferraro (2019), the authors construct a convergence field hybrid quadratic estimator for a Gaussian CMB source field. The hybrid estimator is comprised of an inverse-variance weighted linear combination of a magnification estimator and a shear estimator. It used the shear-only component for small scales and both the convergence and shear components for large scales. The authors found that the hybrid estimator outperformed the SNR of the standard quadratic estimator and that the shear component was less affected by foreground contamination and non-Gaussianities (Schaan & Ferraro, 2019).
@@
Lu & Pen (2008) present a similar hybrid estimator for a 21 cm source field. The authors combined the magnification and shear estimators at all spatial scales. They show that, for a Gaussian source field, their estimator is equivalent to the 3D optimal quadratic deflection estimator introduced by Zahn & Zaldarriaga (2006), but was less affected by gravitational nonlinearities when reconstructions were performed using a non-Gaussian source field. Like Schaan & Ferraro (2019), the authors found that the hybrid estimators were also less susceptible to foreground contamination.
@@ 2.2 Lensing reconstruction with quadratic estimators
The lensing maps are reconstructed using two combinations of the background galaxy density field along with a Fourier space lensing kernel $g_{{\bm{{L}}},{\bm{{\ell}}}}$ that captures the lensing-related correlations. We use a notation where ${\bm{{L}}}$ refers to an angular mode in the foreground lensing potential, while ${\bm{{\ell}}}$ refers to angular modes in the background source that is being lensed.
@@ 2.3 The angular power spectra of LSST galaxy subsamples
Refer to caption (a) $\frac{d{n}}{d{z}}$ Refer to caption (b) $b(z)$ Refer to caption (c) $C_{\ell}$
@@ 2.4 The theoretical convergence power spectrum
Lensing reconstruction using galaxy density fields as a source has great potential due to the massive amounts of galaxy position data available from existing surveys such as DES. This data has not been used for lensing reconstruction, so an efficient reconstruction method could be used to extract information from this otherwise untapped data source. This algorithm could also be used with new surveys such as LSST which provide unprecedented depth and precision in galaxy position measurements (Ivezić et al., 2019; Connolly et al., 2014; Korytov et al., 2019). This information could be used to complement CMB lensing measurements. As an independent measurement, this would help improve the statistical power of observations and identify systematic errors.
@@ 3.1 Convergence power spectra
This is equivalent to Eq. (4.1) in (Nistane et al., 2022). For the survey parameters we use, the signal is much smaller than the noise for all $L$ as seen in Figure 2. Because of this, the SNR of the autocorrelation function will be small. We find that, for the LSST-like survey we use, the total SNR of the autocorrelation function for the standard quadratic estimator is $\left(\frac{S}{N}\right)_{a}=5.25$ when performing integration to $\ell=70,\!000$. Including small scales in the source field has a large effect. If we cut off our integration at $\ell=3000$, the total SNR of the autocorrelation function for the standard quadratic estimator is only 0.52. This is also substantially (two orders of magnitude) smaller than the total signal to noise found in Nistane et al. (2022), as expected, where most of the signal to noise was coming from the inclusion of magnification bias.
@@ 4 Conclusion
A substantial concern that we have neglected for this analysis is the non-Gaussianity of the galaxy number density counts on small scales, exactly the scales where we are obtaining most of our lensing information. Correcting for non-Gaussianities would help further improve our estimator. The non-Gaussianities and strong nonlinearities at small scales in galaxy density fields may make higher-order moments of the galaxy density field (the bispectrum, trispectrum, etc.) non-negligible and introduce additional noise. Line intensity mapping (LIM) source fields are highly nonlinear because they, like the galaxy number counts, trace the underlying matter distribution. Reconstruction methods designed for 21 cm LIM (Pen, 2004; Zahn & Zaldarriaga, 2006; Foreman et al., 2018; Lu et al., 2010) may be applicable for reconstruction using a galaxy density field as a source. Future work should evaluate the impact of non-Gaussianities and foreground contamination on the SNR of reconstructions using our estimator. We would evaluate if, like the estimators described by Lu & Pen (2008) and Schaan & Ferraro (2019), the shear estimator is less susceptible to foreground contamination and non-Gaussianities than the standard quadratic estimator when using a galaxy density source field. The impact of non-Gaussianities should be compared with reconstructions using the methods of Foreman et al. (2018) and Lu et al. (2010) for a galaxy density field background because these estimators are designed to reduce non-Gaussian noise (or, for the case of Lu et al. (2010), are optimal for non-Gaussian fields).

### v2

@@ 1 Introduction
There is large scale structure (LSS) in the universe, as evidenced by the spatial clustering of galaxies (Peebles, 1980). This LSS deflects light as it propagates through the inhomogeneous universe (Bartelmann & Schneider, 2001). These deflections can occasionally lead to spectacular Einstein rings and multiple images, but usually will lead to only small distortions to the shapes and brightnesses of background objects. This has been measured in galaxy surveys through the changes in the detailed shapes of galaxies (Luo et al., 2018; Gatti et al., 2021), and has been measured in the distortion to the characteristic clustering patterns in the cosmic microwave background (Aghanim et al., 2020; Omori et al., 2023; Pan et al., 2023; Madhavacheril et al., 2023; Qu et al., 2024; Ade et al., 2014; BICEP/Keck Collaboration et al., 2022).
@@
In the case of CMB lensing, the workhorse tool that is widely used is a quadratic estimator (Okamoto & Hu, 2003), which amounts to multiplying together two CMB maps that each have their own filtering. The filtering of the maps is chosen to create an unbiased, minimum variance estimator, which optimizes the SNR of the resulting lens reconstruction. If this lens reconstruction is then correlated with a galaxy-based LSS tracer, the result could be expressed as a 3-pt function, as it involves products of two CMB maps and an LSS map. As we will show below, for the similar case of lensing of the galaxy clustering distribution by foreground matter, the reconstruction of the foreground lensing potential will involve two versions of the background galaxy number count field, and then cross-correlation with an LSS tracer such as cosmic shear can be expressed as a galaxy-galaxy-shear 3-pt function. If we instead used the foreground galaxy counts, then there is a 3-pt function that would be quadratic in the background galaxy number counts and linear in foreground galaxy number counts. In general, the three fields going into these estimates can be from diverse LSS tracers: the background 2-pt function could be cross-correlations between different LSS tracers, and the foreground could be taken to be any LSS tracer of the foreground matter (Darwish et al., 2021). In this work, we only look at the simplest case of using the auto-correlation of galaxies in the background being lensed by foreground matter in a way that leads to a measured shear of the same background material, mainly to keep the accounting simple.
@@
An important point of focus in each of these studies has been the non-Gaussianity of the background field being lensed. It has been shown (Schaan et al., 2018; Foreman et al., 2018; Feng & Holder, 2019) that there can be substantial biases in the recovered lensing power spectrum if the non-Gaussianity is ignored. Nevertheless, for this preliminary investigation we will disregard the non-Gaussianity of the field and its uncertainties in what follows, so the results presented below will no doubt require future work to study these effects. These corrections arise because non-linear gravitational collapse leads to an intrinsic background of spatial variation in the statistics of small-scale clustering, the magnitude of which depends on factors including the spatial scales probed, redshift, and survey parameters (Foreman et al., 2018). However, some lensing estimators that are optimal for a Gaussian source field have been shown to be resistant to non-Gaussian noise. Foreman et al. (2018) constructed a bias-hardened estimator that included contributions from the density field trispectrum that corrected for mild gravitational nonlinearities, while Lu et al. (2010) calibrated their estimator with N-body simulations to create a set of lensing weights that are optimal for non-Gaussian fields.
@@
In this work, we closely follow the procedures described by Schaan & Ferraro (2019), who present a method for lensing reconstruction that is less affected by foreground contamination and non-Gaussianities. In Schaan & Ferraro (2019), the authors construct a convergence field hybrid quadratic estimator for a Gaussian CMB source field. The hybrid estimator is comprised of an inverse-variance weighted linear combination of a magnification estimator and a shear estimator. It used the shear-only component for small scales and both the convergence and shear components for large scales. The authors found that the hybrid estimator outperformed the SNR of the standard quadratic estimator and that the shear component was less affected by foreground contamination (Schaan & Ferraro, 2019). Qu et al. (2023) generalized this estimator for full-sky reconstruction and verified using simulations that these estimators are less affected by foreground contamination than the standard QE.
The way that foregrounds contaminate CMB lensing measurements is through their non-Gaussianity; reducing foreground contamination means reducing the impact of the non-Gaussianity induced by non-linear structure formation. It is therefore expected that these new estimators should also be less sensitive to gravitationally induced non-Gaussianity of the background source field. Schaan et al. (2018) conjectured that a shear-only estimator would be robust to non-Gaussian contamination for CIB fields, which consist of a superposition of point sources (galaxies) that trace the nonlinear density field. Furthermore, in what follows below we will primarily focus on cross-correlation, where the impact of source field non-Gaussianity may also be lower.
@@
Lu & Pen (2008) present a similar hybrid estimator for a 21 cm source field. The authors combined the magnification and shear estimators at all spatial scales. They show that, for a Gaussian source field, their estimator is equivalent to the 3D optimal quadratic deflection estimator introduced by Zahn & Zaldarriaga (2006), but was less affected by gravitational nonlinearities when reconstructions were performed using a non-Gaussian source field. Like Schaan & Ferraro (2019), the authors found that the hybrid estimators were also less susceptible to this foreground contamination.
The estimator we use (Schaan & Ferraro, 2019) was developed for CMB lensing reconstruction, which uses a Gaussian source field. We propose here to perform reconstructions using a galaxy number count source field, which is highly non-Gaussian, and detailed simulations will ultimately be required to assess the robustness of the shear estimator for this application. The 21 cm source field, with large fluctuations in ionization in addition to the density fluctuations, is expected to be substantially more non-Gaussian than a typical galaxy distribution, so the results of Lu & Pen (2008) are promising in finding that the shear estimator is not substantially biased by this non-Gaussianity.
Typical CMB lensing techniques probe large spatial scales ($\ell\lessapprox 3,\!000$), but we found that most of the SNR for the cross-correlation between the observed and theoretical power spectra is contained at very large $\ell$ (up to $\ell\approx 70,\!000$). These are similar scales to what was explored in both Lu & Pen (2008) and Schaan et al. (2018), where there were no obvious problems with applying these techniques other than the non-Gaussianity of the field leading to the traditional weights becoming sub-optimal.
@@ 2.2 Lensing reconstruction with quadratic estimators
The lensing maps are reconstructed using a quadratic estimator that uses two combinations of the background galaxy density field along with a Fourier space lensing kernel $g_{{\bm{{L}}},{\bm{{\ell}}}}$ that captures the lensing-related correlations. We use a notation where ${\bm{{L}}}$ refers to an angular mode in the foreground lensing potential, while ${\bm{{\ell}}}$ refers to angular modes in the background source that is being lensed.
@@ 2.3 The angular power spectra of LSST galaxy subsamples
(a) $\frac{d{n}}{d{z}}$
(b) $b(z)$
(c) $C_{\ell}$
@@ 2.4 The theoretical convergence power spectrum
Lensing reconstruction using galaxy density fields as a source has great potential due to the massive amounts of galaxy position data available from existing surveys such as DES. This data has not been used for lensing reconstruction, so an efficient reconstruction method could be used to extract information from this otherwise untapped data source. This algorithm could also be used with new surveys such as LSST which provide unprecedented depth and precision in galaxy position measurements (Ivezić et al., 2019; Connolly et al., 2014; Korytov et al., 2019). This information could be used to complement CMB lensing measurements in the same way that cosmic shear measurements can be used. This is an independent measurement of the same physical effect as traditional galaxy shear measurements and is therefore a good complement to cosmic shear measurements. This could in principle help improve the statistical power of observations and identify possible systematic errors such as multiplicative errors in shear measurements. We could use the same galaxies for the shear measurements as for the lensing reconstruction, in which case the galaxy shape distortions and the distortion of the clustering should be a direct test for systematic errors.
@@ 3.1 Convergence power spectra
This is equivalent to Eq. (4.1) in (Nistane et al., 2022). For the survey parameters we use, the signal is much smaller than the noise for all $L$ as seen in Figure 2. Because of this, the SNR of the autocorrelation function will be small. We find that, for the LSST-like survey we use, the total SNR of the autocorrelation function for the standard quadratic estimator is $\left(\frac{S}{N}\right)_{a}=5$ when performing integration to $\ell=70,\!000$. We sum over all $L$, but find that 75% of the SNR comes from $L<100$ and 98.5% comes from $L<3,\!000$. While small scales in the lensing field contribute little to the SNR of the autocorrelation spectrum because noise dominates this region, including small scales from the source field has a large effect, a result that is consistent with Lu & Pen (2008) and Schaan et al. (2018). If we cut off our integration at $\ell=3,\!000$, the total SNR of the autocorrelation function for the standard quadratic estimator is only 0.5. This is also substantially (two orders of magnitude) smaller than the total signal to noise found in Nistane et al. (2022), as expected, where most of the signal to noise was coming from the inclusion of magnification bias.
@@ 4 Conclusion
For this forecast, we calculate the galaxy power spectrum using halofit and the linear bias, which are not accurate at very small scales. In practice, the galaxy power spectrum will be already well-measured before performing the lensing reconstruction. This introduces uncertainty in our forecast, but uncertainty in the galaxy power spectrum will not be a limiting factor for data analysis. This uncertainty in the accuracy of the non-linear matter power spectrum calculation will be important for interpreting the reconstructed convergence spectra because some of the SNR comes from small scales, especially in the limit of a low-noise tracer of the density field being used for cross-correlation.
A substantial concern that we have neglected for this analysis is the non-Gaussianity of the galaxy number density counts on small scales, exactly the scales where we are obtaining most of our lensing information. Correcting for non-Gaussianities would help further improve our estimator. The non-Gaussianities and strong nonlinearities at small scales in galaxy density fields may make higher-order moments of the galaxy density field (the bispectrum, trispectrum, etc.) non-negligible and introduce additional noise. Line intensity mapping (LIM) source fields are highly nonlinear because they, like the galaxy number counts, trace the underlying matter distribution. Reconstruction methods designed for 21 cm LIM (Pen, 2004; Zahn & Zaldarriaga, 2006; Foreman et al., 2018; Lu et al., 2010) may be applicable for reconstruction using a galaxy density field as a source. Future work should evaluate the impact of non-Gaussianities and foreground contamination on the SNR of reconstructions using our estimator. We would evaluate if, like the estimators described by Lu & Pen (2008) and Schaan et al. (2018), the shear estimator is less susceptible to foreground contamination than the standard quadratic estimator when using a galaxy density source field. The impact of non-Gaussianities should be compared with reconstructions using the methods of Foreman et al. (2018) and Lu et al. (2010) for a galaxy density field background because these estimators are designed to reduce non-Gaussian noise (or, for the case of Lu et al. (2010), are optimal for non-Gaussian fields). The impact of non-Gaussianities, as well as the robustness of the QE when performing $\ell$-integration to scales much smaller than are typically used for CMB lensing reconstruction, ultimately must be evaluated by performing lensing reconstructions on mock catalogs, a task left for future work.